OpenStudy (anonymous):

An advertising balloon has radius 6m. Air leak out at a rate of 28 pi m^3/minute. Find the rate of decrease of the diameter after 9 minute ?

OpenStudy (anonymous):

what is the volume of a sphere in term of the radius?

OpenStudy (anonymous):

oh i think it is \[V=\frac{4}{3}\pi r^3\]

OpenStudy (anonymous):

taking derivatives you get \[V'=4\pi r^2 r'\] you are told \(V'=28\) solve for \(r'\)

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

how about 9 min ?

OpenStudy (anonymous):

i made an equation like this: 28= 4/3π3r dr/dt

OpenStudy (anonymous):

i mean r^2

OpenStudy (anonymous):

yes, and you can find \(r\) because you know that initially it was 6 m and you know the rate of decrease

OpenStudy (anonymous):

thank you :)